. If you like to buy a new equipment for $13000 and you receive a cash inflow of $ 2500 per year for 10 years. What is the internal rate of return?



1
Expert's answer
2021-12-05T18:56:21-0500

Net Present Value:

"NPV=\sum\frac{CF_t}{(1+r)^t}=13000"

where CFt is cash inflow during a single period t,

r is internal rate of return


so, we have:

"\sum\frac{2500}{(1+r)^t}=13000"

t= 10 years

using online calculator https://www.calculatestuff.com/financial/irr-calculator, we get:

"r=14.08\%"